Fred and Kate are brother and sister. Fred has three times as many sisters as brothers, and Kate has as many brothers as sisters. How many boys and girls are there in this family?
Answer by ankor@dixie-net.com(22740)

Fred and Kate are brother and sister.
Fred has three times as many sisters as brothers, and Kate has as many brothers as sisters.
How many boys and girls are there in this family?
:
Let b = no. of boys
let b = no. of girls
:
Fred:
g = 3(b-1)
Kate
g-1 = b
g = b + 1
replace g with (b+1) in the first equation
b + 1 = 3b - 3
1 + 3 = 3b - b
4 = 2b
b = 2 boys
then
g = 2 + 1
g = 3 girls